OpenStudy (anonymous):

integrate cos x = sin x.......ok integrate cos 2x = what is my formula here? integrate cos 3x...isn't there some formula for all of these?

OpenStudy (experimentx):

there is a technique called substitution ... let \( 2x = u\) then \( 2 dx = du \) or \( dx = du/2\) Then, \[ \int \cos2x\;dx = \int \cos u \;du/2 = 1/2 * \sin u = 1/2 * \sin 2x\]

OpenStudy (experimentx):

Try these substitution for others, you will get general formula!! \[ \int \cos nx \;dx = \frac{\sin nx}{n} + c\]

OpenStudy (anonymous):

ok think i am there ---was putting in the question wrong over at wolfram, remember square brackets after the cos or sin or -cos or -sin